K= 1.44 at 1000 K for synthesis gas reaction:

CO(g)+H2O←→CO2(g)+H2(g)

If initially the gases are introduced into an empty vessel at a partial pressure at 1.0 ATM each,what will be the equilibrium pressure of each gas?


1
Expert's answer
2021-06-14T13:38:20-0400

Total moles = 4

Partial pressure = 0.25(1)

= 0.25 for each

Kp = 1/1/1/1 = 1

O.25(1)= 4X2

0.25 -4x2 = 0

0.25 - 4x2-0

X =- b+/ - (✓b² - 4ac)/2a

= 4 +/-( √b2 - 0)/0

= 4 +/- 0

= 4

x = 4

= 4 atm